Cost of Units Transferred Out and Ending Work in Process

The costs per equivalent unit of direct materials and conversion in the Rolling Department of Kraus Steel Company are $750 and $120, respectively. The equivalent units to be assigned costs are as follows:

Equivalent Units
Direct Materials Conversion
Inventory in process, October 1 0 80
Started and completed during October 3,700 3,700
Transferred out of Rolling (completed) 3,700 3,780
Inventory in process, October 31 300 75
Total units to be assigned costs 4,000 3,855

The beginning work in process inventory on October 1 had a cost of $163,800. Determine the cost of completed and transferred-out production, the ending work in process inventory, and the total costs assigned by the Rolling Department.

Completed and transferred-out production $
Inventory in process, October 31 $
Total costs assigned by the Rolling Department $
Answer #1

Completed and transferred out production = (3700*750+3780*120+163800) = 3392400

Inventory in process, October 31 = (300*750+75*120) = 234000

Total Cost assigned by the rolling department = 234000+3392400 = 3626400